Luar1118 auuvily 2. Ivivuuiu v Name Date Show all work or no credit will be given. 1. The height of a ball (in meters) aboveht=-4.9t2+4.2t this function
Answer #1

1. A.  The height of a ball above the ground ( in meters) after t seconds is given by h(t) = -4.9t2+4.2t = -4.9[t2- (4.2/4.9)t] = -4.9(t2- 6t/7) = -4.9[t2- 2t *3/7 + (3/7)2] +4.9*(3/7)2= -4.9(t-3/7)2 + 9/10.

This is the vertex form of the equation of a downwards opening parabola with vertex at (3/7,9/10).

B. A graph of the function h(t) = -4.9t2+4.2t is attached. Since the vertex of a downwards opening parabola is its highest point, hence the y -coordinate of the vertex i.e. 9/10 meters is the maximum height attained by the ball. This height is attained 3/7 second after the ball is thrown up.